# DOCE: Finding the Sweet Spot for Execution-Based Code Generation

### v1

@@ 6 Conclusion and Future Work
We leverage recent advances in code generation, including N-best reranking, MBR, and self-debugging and we propose DOCE, a unified framework for generating code with higher execution accuracy. We perform an extensive study with our proposed framework on reranking by studying candidate generation, comparing across reranking methods, and studying self-debugging for increasing reranking upper bound. Importantly, we highlight the impact of filtering based on trial unit tests, a commonly used technique whose effect has been overlooked in previous works, showing that it is crucial for all procedures DOCE. Last but not least, we observe that self-debugging improves the upper bound, with self-debugging on all generated candidates giving further improvement of the reranking results compared to self-debugging on one single candidate. Potential future works include incorporating studying the effects of generated unit tests and specifications, studying the effect of directly and indirectly using feedback from trial unit tests, and training better general-purpose execution-less metrics for reranking.
@@
First of all, our paper is limited by the scale of experiments on more code generation models, due to the limit of compute and the limit. Our solution for this is to select representative classes of open-source models and experiments on models with parameters below 15 billions.
Appendix A Full Results
A.1 Candidate Generation
We present the effect of the number of candidates with candidates generated by CodeLlama-{7,13}B-Instruct (see Figure 8 and Figure 10) DeepSeekCoder-6.7B-Instruct (see Figure 9) on both the original unit tests and extended test cases in EvalPlus (Liu et al., 2023). Most results align with those we present in Figure 3, where filtering based on trial unit tests improves MBR-Exec considerably and approaches the oracle performance when combined with MBR-Exec. The only exception happens on candidates generated by DeepSeekCoder-6.7B-Instruct on HumanEval (see Figure 9(a)) where the execution accuracy using MBR-Exec with filtering on no less than 10 candidates is already above 90.
@@ A.1 Candidate Generation
A.2 Comparing Reranking Methods
@@
A.3 MBR-Exec with Fewer Unit Tests
@@
A.4 Improving Oracle with Self-Debugging
@@
We first present results of improvement of Pass@k that estimates oracle improvement using candidates generated and self-debugged by CodeLlama-7B-Instruct (see Figure 14), DeepSeekCoder-6.7B-Instruct (see Table 15) and CodeLlama-13B-Instruct (see Table 16). Our findings align with Section 4.4 as one round of self-debugging is enough to improve the oracle.
@@
We then present results over different numbers of candidates generated and debugged by CodeLlama-7B-Instruct (see Figure 17), DeepSeekCoder-6.7B-Instruct (see Table 18) and CodeLlama-13B-Instruct (see Table 19). We find that SD-Multi outperforms SD-1 constantly except for HumanEval(+) with candidates generated and debugged by DeepSeekCoder-6.7B-Instruct with 25 generated candidates.
Appendix B Prompts
B.1 Prompt for Code Generation (0-Shot)
1[INST] Write code to solve the following coding problem that obeys the constraints and passes the example test cases. Please wrap your code answer using ‘‘‘:
2‘‘‘python
3{prompt}
4‘‘‘
5[/INST]
6‘‘‘python
Figure 20: 0-shot prompt for code generation with CodeLlama-{7,13}B-Instruct.
1You are an AI programming assistant, utilizing the DeepSeek Coder model, developed by DeepSeek Company, and you only answer questions related to computer science. For politically sensitive questions, security and privacy issues, and other non-computer science questions, you will refuse to answer.
2### Instruction:
3Please complete the following Python function in a markdown-style code block:
4‘‘‘python
5{prompt}
6‘‘‘
7### Response:
8‘‘‘python
Figure 21: 0-shot prompt for code generation with DeepSeekCoder-6.7B-Instruct.
See Figure 20 and Figure 21.
B.2 Prompt for Reviewer (3-Shot)
1### task start ###
2### code
3‘‘‘python
4def count_ways(n):
5 A = [0] * (n + 1)
6 B = [0] * (n + 1)
7 A[0] = 1
8 A[1] = 0
9 B[0] = 0
10 B[1] = 1
11 for i in range(2, n+1):
12 A[i] = A[i - 2] + 2 * B[i - 1]
13 B[i] = A[i - 1] + B[i - 2]
14 return A[n]
15‘‘‘
16
17Write the docstring for the above code.
18
19### docstring
20‘‘‘python
21"""
22Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
23$>$ count_ways(2)
243
25"""
26‘‘‘
27### task end ###
28
29### task start ###
30### code
31‘‘‘python
32import re
33def find_char_long(text):
[…]
35‘‘‘
36
37Write the docstring for the above code.
38
39### docstring
40‘‘‘python
41"""
42Write a function to find all words which are at least 4 characters long in a string by using regex.
43$>$ find_char_long(’Please move back to stream’)
44[’Please’, ’move’, ’back’, ’stream’]
45"""
46‘‘‘
47### task end ###
48
49### task start ###
50### code
51‘‘‘python
52def square_nums(nums):
53 square_nums = list(map(lambda x: x ** 2, nums))
54 return square_nums
55‘‘‘
56
57Write the docstring for the above code.
58
59### docstring
60‘‘‘python
61"""
62Write a function to find squares of individual elements in a list using lambda function.
63$>$ square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])
64[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
65"""
66‘‘‘
67### task end ###
1### task start ###
2### code
3‘‘‘python
4{code}
5‘‘‘
6
7Write the docstring for the above code.
8
9### docstring
10‘‘‘python
11{prompt}
12‘‘‘
Figure 22: 3-shot prompt for getting Reviewer log-likelihoods (Zhang et al., 2022) on HumanEval.
1### task start ###
2### code
3‘‘‘python
4def count_ways(n):
5 A = [0] * (n + 1)
6 B = [0] * (n + 1)
7 A[0] = 1
8 A[1] = 0
9 B[0] = 0
10 B[1] = 1
11 for i in range(2, n+1):
12 A[i] = A[i - 2] + 2 * B[i - 1]
13 B[i] = A[i - 1] + B[i - 2]
14 return A[n]
15‘‘‘
16
17Write the docstring for the above code.
18
19### docstring
20‘‘‘python
21"""
22Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
23assert count_ways(2) == 3
24"""
25‘‘‘
26### task end ###
27
28### task start ###
29### code
30‘‘‘python
31import re
32def find_char_long(text):
[…]
34‘‘‘
35
36Write the docstring for the above code.
37
38### docstring
39‘‘‘python
40"""
41Write a function to find all words which are at least 4 characters long in a string by using regex.
42assert find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]
43"""
44‘‘‘
45### task end ###
1### task start ###
2### code
3‘‘‘python
4def square_nums(nums):
5 square_nums = list(map(lambda x: x ** 2, nums))
6 return square_nums
7‘‘‘
8
9Write the docstring for the above code.
10
11### docstring
12‘‘‘python
13"""
14Write a function to find squares of individual elements in a list using lambda function.
15assert square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])==[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
16"""
17‘‘‘
18### task end ###
19
20### task start ###
21### code
22‘‘‘python
23{code}
24‘‘‘
25
26Write the docstring for the above code.
27
28### docstring
29‘‘‘python
30{prompt}
31‘‘‘
Figure 23: 3-shot prompt for getting Reviewer log-likelihoods (Zhang et al., 2022) on MBPP.
See Figure 22 and Figure 23.
B.3 Prompt for Self-Debugging (6-Shot)
1### task start ###
2### prompt
3‘‘‘python
4def count_ways(n):
5 """ Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
6 $>$ count_ways(2)
7 3
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def count_ways(n):
14 if n == 0:
15 return 1
16 if n == 1:
17 return 1
18 if n == 2:
19 return 3
20 return count_ways(n-1) + count_ways(n-2)
21‘‘‘
22
23### feedback
24With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
25The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
26
27### fixed code
28‘‘‘python
29def count_ways(n):
30 A = [0] * (n + 1)
31 B = [0] * (n + 1)
32 A[0] = 1
33 A[1] = 0
34 B[0] = 0
35 B[1] = 1
36 for i in range(2, n+1):
37 A[i] = A[i - 2] + 2 * B[i - 1]
38 B[i] = A[i - 1] + B[i - 2]
39 return A[n]
40‘‘‘
41
42### feedback
43With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
44The code above is correct.
45### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def differ_At_One_Bit_Pos(lhs,rhs):
5 """ Write a python function to check whether the two numbers differ at one bit position only or not.
6 $>$ differ_At_One_Bit_Pos(15,8)
7 False
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def differ_At_One_Bit_Pos(lhs,rhs):
14 """ Write a python function to check whether the two numbers differ at one bit position only or not.
15 $>$ differ_At_One_Bit_Pos(15,8)
16 False
17 """
18 if (lhs - rhs) == 0 or (lhs - rhs) == 1:
19 return True
20 return False
21‘‘‘
22
23### feedback
24With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
25The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
26
27### fixed code
28‘‘‘python
29def differ_At_One_Bit_Pos(a,b):
30 """ Write a python function to check whether the two numbers differ at one bit position only or not.
31 $>$ differ_At_One_Bit_Pos(15,8)
32 False
33 """
34 x = a ^ b
35 return x and (not(x & (x - 1)))
36‘‘‘
37
38### feedback
39With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
40The code above is correct.
41### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def find_char_long(text):
5 """ Write a function to find all words which are at least 4 characters long in a string by using regex.
6 $>$ find_char_long(’Please move back to stream’)
7 [’Please’, ’move’, ’back’, ’stream’]
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def find_char_long(text):
[…]
15‘‘‘
16
17### feedback
18With the above function, theres following error:
19"""
20NameError: name ’re’ is not defined
21"""
22The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
23
24### fixed code
25‘‘‘python
26import re
27def find_char_long(text):
[…]
29‘‘‘
30
31### feedback
32With the above function, the assertion is find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]‘ and the real execution output is ‘[’Please’, ’move’, ’back’, ’stream’]‘. The code passes the assertion.
33The code above is correct.
34### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def square_nums(nums):
5 """ Write a function to find squares of individual elements in a list using lambda function.
6 $>$ square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])
7 [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def square_nums(nums):
14 """ Write a function to find squares of individual elements in a list using lambda function.
15 $>$ square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])
16 [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
17 """
18 square_nums = list(map(lambda x: x ** 2, nums))
19 return square_nums
20‘‘‘
21
22### feedback
23With the above function, the assertion is square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10]) == [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘ and the real execution output is ‘[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘. The code passes the assertion.
24The code above is correct.
25### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def find_Rotations(s):
5 """ Write a python function to find the minimum number of rotations required to get the same string.
6 $>$ find_Rotations("aaaa")
7 1
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def find_Rotations(s):
14 if len(s) == 1:
15 return 1
16 elif len(s) == 0:
17 return 0
18 else:
19 l = len(s)
20 min = len(s)
21 for i, c in enumerate(s):
22 if c == s[0]:
23 temp = i
24 if temp < min:
25 min = temp
26 return min
27‘‘‘
28
29### feedback
30With the above function, the assertion is find_Rotations("aaaa") == 1‘ but the real execution output is ‘0‘.
31The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
32
33### fixed code
34‘‘‘python
35def find_Rotations(s):
36 n = len(s)
37 for i in range(1, n + 1):
38 if s[i:n] + s[0:i] == s:
39 return i
40 return n
41‘‘‘
42
43### feedback
44With the above function, the assertion is find_Rotations("aaaa") == 1‘ and the real execution output is ‘1‘. The code passes the assertion.
45The code above is correct.
46### task end ###
1### task start ###
2### prompt
3‘‘‘python
4def small_nnum(lst,n):
5 """ Write a function to get the n smallest items from a dataset.
6 $>$ small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2)
7 [10, 20]
8 """
9‘‘‘
10
11### code
12‘‘‘python
13def small_nnum(lst,n):
14 """ Write a function to get the n smallest items from a dataset.
15 $>$ small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2)
16 [10, 20]
17 """
18 lst = sorted(lst)
19 lst = lst[:n]
20 return lst
21‘‘‘
22
23### feedback
24With the above function, the assertion is small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2) == [10,20]‘ and the real execution output is ‘[10, 20]‘. The code passes the assertion.
25The code above is correct.
26### task end ###
1### task start ###
2### instruction
3{instruction}
4‘‘‘python
5{unit_test}
6‘‘‘
7
8### code
9‘‘‘python
10{code}
11‘‘‘
12
13###feedback
14{feedback}
15The code above is
Figure 24: 6-shot prompt for self-debugging on HumanEval.
1### task start ###
2### instruction
3Write a function to find the number of ways to fill it with 2 x 1 dominoes for the given 3 x n board.
4‘‘‘python
5assert count_ways(2) == 3
6‘‘‘
7
8### code
9‘‘‘python
10def count_ways(n):
11 if n == 0:
12 return 1
13 if n == 1:
14 return 1
15 if n == 2:
16 return 3
17 return count_ways(n-1) + count_ways(n-2)
18‘‘‘
19
20### feedback
21With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
22The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
23
24### fixed code
25‘‘‘python
26def count_ways(n):
27 A = [0] * (n + 1)
28 B = [0] * (n + 1)
29 A[0] = 1
30 A[1] = 0
31 B[0] = 0
32 B[1] = 1
33 for i in range(2, n+1):
34 A[i] = A[i - 2] + 2 * B[i - 1]
35 B[i] = A[i - 1] + B[i - 2]
36 return A[n]
37‘‘‘
38
39### feedback
40With the above function, the assertion is count_ways(2) == 3‘ and the real execution output is 3. The code passes the assertion.
41The code above is correct.
42### task end ###
1### task start ###
2### instruction
3Write a python function to check whether the two numbers differ at one bit position only or not.
4‘‘‘python
5assert differ_At_One_Bit_Pos(15,8) == False
6‘‘‘
7
8### code
9‘‘‘python
10def differ_At_One_Bit_Pos(lhs,rhs):
11 if (lhs - rhs) == 0 or (lhs - rhs) == 1:
12 return True
13 return False
14‘‘‘
15
16### feedback
17With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
18The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
19
20### fixed code
21‘‘‘python
22def differ_At_One_Bit_Pos(a,b):
23 x = a ^ b
24 return x and (not(x & (x - 1)))
25‘‘‘
26
27### feedback
28With the above function, the assertion is differ_At_One_Bit_Pos(15,8) == False and the real execution output is False‘. The code passes the assertion.
29The code above is correct.
30### task end ###
1### task start ###
2### instruction
3Write a function to find all words which are at least 4 characters long in a string by using regex.
4‘‘‘python
5assert find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]
6‘‘‘
7
8### code
9‘‘‘python
10def find_char_long(text):
[…]
12‘‘‘
13
14### feedback
15With the above function, theres following error:
16’’
17NameError: name re is not defined
18’’
19The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
20
21### fixed code
22‘‘‘python
23import re
24def find_char_long(text):
[…]
26‘‘‘
27
28### feedback
29With the above function, the assertion is find_char_long(’Please move back to stream’) == [’Please’, ’move’, ’back’, ’stream’]‘ and the real execution output is ‘[’Please’, ’move’, ’back’, ’stream’]‘. The code passes the assertion.
30The code above is correct.
31### task end ###
1### task start ###
2### instruction
3Write a function to find squares of individual elements in a list using lambda function.
4‘‘‘python
5assert square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10])==[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]
6‘‘‘
7
8### code
9‘‘‘python
10def square_nums(nums):
11 square_nums = list(map(lambda x: x ** 2, nums))
12 return square_nums
13‘‘‘
14
15### feedback
16With the above function, the assertion is square_nums([1, 2, 3, 4, 5, 6, 7, 8, 9, 10]) == [1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘ and the real execution output is ‘[1, 4, 9, 16, 25, 36, 49, 64, 81, 100]‘. The code passes the assertion.
17The code above is correct.
18### task end ###
1### task start ###
2### instruction
3Write a python function to find the minimum number of rotations required to get the same string.
4‘‘‘python
5assert find_Rotations("aaaa") == 1
6‘‘‘
7
8### code
9‘‘‘python
10def find_Rotations(s):
11 if len(s) == 1:
12 return 1
13 elif len(s) == 0:
14 return 0
15 else:
16 l = len(s)
17 min = len(s)
18 for i, c in enumerate(s):
19 if c == s[0]:
20 temp = i
21 if temp < min:
22 min = temp
23 return min
24‘‘‘
25
26### feedback
27With the above function, the assertion is find_Rotations("aaaa") == 1‘ but the real execution output is ‘0‘.
28The code above is incorrect. Please fix it. Make sure the code included in instruction will be executed correctly. Watch out for function/variable names, exception handling, and other errors.
29
30### fixed code
31‘‘‘python
32def find_Rotations(s):
33 n = len(s)
34 for i in range(1, n + 1):
35 if s[i:n] + s[0:i] == s:
36 return i
37 return n
38‘‘‘
39
40### feedback
41With the above function, the assertion is find_Rotations("aaaa") == 1‘ and the real execution output is ‘1‘. The code passes the assertion.
42The code above is correct.
43### task end ###
1### task start ###
2### instruction
3Write a function to get the n smallest items from a dataset.
4‘‘‘python
5assert small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2) == [10,20]
6‘‘‘
7
8### code
9‘‘‘python
10def small_nnum(lst,n):
11 lst = sorted(lst)
12 lst = lst[:n]
13 return lst
14‘‘‘
15
16### feedback
17With the above function, the assertion is small_nnum([10, 20, 50, 70, 90, 20, 50, 40, 60, 80, 100],2) == [10,20]‘ and the real execution output is ‘[10, 20]‘. The code passes the assertion.
18The code above is correct.
19### task end ###
1### task start ###
2### instruction
3{instruction}
4‘‘‘python
5{unit_test}
6‘‘‘
7
8### code
9‘‘‘python
10{code}
11‘‘‘
12
13###feedback
14{feedback}
15The code above is
Figure 25: 6-shot prompt for self-debugging on MBPP-S.
See Figure 24 and Figure 25.

### v4

@@ 6 Conclusion and Future Work
We propose DOCE, a unified inference framework for code generation. We systematically study candidate generation, reranking, and self-debugging. Importantly, we highlight the impact of high-temperature sampling, execution-based reranking with few high-quality unit tests, and self-debugging with multiple candidates. Potential future works include analyzing unit test quality for reranking and proposing better methods for generating code candidates, exploring the current framework with generated unit test cases, and exploring general-purpose rerankers.
@@
First of all, our paper is limited by the scale of experiments as we cannot experiment with all available large language models due to the limit of computation, and our solution is to select representative classes of open-source models and experiments on models with parameters at most 16 billions. Additionally, this paper is limited by potential risks of data contamination from the dataset, despite that HumanEval and MBPP are commonly used in evaluating code generation models. Our solution is to include experiments with LiveCodeBench, which contains a lower risk of data contamination.
Ethical Considerations
We do not consider the existence of ethical issues related to the paper, due to the nature of code generation and our usage of publicly available datasets that have been verified. However, we noticed the risk of ethical concern due to our choice of sampling temperature. We checked generations and found no ethical issues in the generated content.
Appendix A Dataset Statistics
We use the HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) datasets in EvalPlus (Liu et al., 2023), consisting of 164 and 395 problems respectively. For LiveCodeBench (Jain et al., 2024), we use the version that includes competitive programming problems from July 2023 to September 2024 to balance the risk of data contamination and size of the dataset. We adopt a simpler evaluation by eliminating all test cases with expected output being extreme values and set 20 unit test cases per problem at most. For easy implementation, we only include problems presented as function completion, obtaining 283 problems, among which 88 are easy problems, 135 are medium-level problems, and 60 are hard problems.
Appendix B Experimental Results
B.1 Candidate Generation
We present the effect of the number of candidates with candidates generated by CodeLlama-{7,13}B-Instruct (see Figure 7 and Figure 9) DeepSeekCoder-{6.7B, V2-Lite}-Instruct (see Figure 8 and Figure 10) on both the original unit tests and extended test cases in EvalPlus (Liu et al., 2023). We also provide all cases for LiveCodeBench (see Figure 11). Most results align with those we present in Figure 2, where filtering based on trial unit tests improves MBR-Exec considerably and approaches the oracle performance when combined with MBR-Exec. The only exception happens on candidates generated by DeepSeekCoder-6.7B-Instruct on HumanEval (see Figure 8(a)) where the execution accuracy using MBR-Exec with filtering on no less than 10 candidates is already above 90.
@@ B.1 Candidate Generation
B.2 Comparing Reranking Methods
@@
B.3 MBR-Exec with Fewer Unit Tests
@@
B.4 Improving Oracle with Self-Debugging
@@
We first present results of improvement of Pass@k that estimates oracle improvement using candidates generated and self-debugged by CodeLlama-7B-Instruct (see Figure 15), DeepSeekCoder-6.7B-Instruct (see Figure 16), CodeLlama-13B-Instruct (see Figure 17), and DeepSeekCoder-V2-Lite-Instruct (see Figure 18). We also provide cases for LiveCodeBench (see Figure 19). Our findings align with Section 4.4 as one round of self-debugging is enough to improve the oracle.
@@
We then present results over different numbers of candidates generated and debugged by CodeLlama-7B-Instruct (see Figure 20), DeepSeekCoder-6.7B-Instruct (see Figure 21), CodeLlama-13B-Instruct (see TFigure 22), and DeepSeekCoder-V2-Lite-Instruct (see Figure 23). We also provide all cases for LiveCodeBench (see Figure 24). We find that SD-Multi outperforms SD-1 constantly except for HumanEval(+) with candidates generated and debugged by DeepSeekCoder-6.7B-Instruct with 25 generated candidates.